# Summarization of Opinionated Political Documents with Varied Perspectives

### v1

@@ 5.2 Extraction Analyses
To further understand model performance on this task, we analyze extraction behavior with respect to different qualities of the input documents: perspective, position in the input, length, and use of arousal terms. We use extractive fragment coverage to quantify this behavior, as defined by Grusky et al. (2018). In these analyses, we hypothesize that models with high performance tend to extract more from the intended perspective. We evaluate three additional hypotheses illustrating potentially undesirable behaviors: models extract more information from documents that (1) appear earlier in the input (position bias), (2) are longer (length bias), and (3) that contain more arousing language (arousal bias).
@@
As the Mixed-Perspective setting requires models to disentangle perspective among input documents, we first analyze whether models tend to extract from source documents with the correct political slant. Figure 3 presents the extraction coverage scores for each combination of summary and source document perspective (e.g., left/right perspective summary and left/right-leaning source documents). For both perspectives, performant models tend to extract from source documents portraying the intended perspective (blue bars in the "Left Summaries" plot, and red bars in "Right Summaries") significantly more often than the alternative perspective. T5, Flan-T5, and Mistral-Instruct, however, present insignificant differences in extraction coverage for both perspectives. Uniquely, BART consistently extracts significantly more from right-leaning documents than left-leaning, regardless of the intended perspective. Therefore, the largest and best performing models (e.g., GPT-4o, Llama-3-70B, etc.) appear to be sufficiently capable of distinguishing perspectives in inputs.
Position and lead biases in summarization systems are well-documented phenomena Chhabra et al. (2024); Olabisi and Agrawal (2024); Jung et al. (2019). Following these works, Figure 4 presents the extraction coverage scores across document positions for each model. As in prior works, we find a near monotonically decreasing relationship between extraction from a document and its position in the input, with all model drawing significantly more terms from the first document compared to those that follow. Among models, Llama-3 (70B) and GPT-4o appear to be most robust to position biases, indicated by the flatter curves, but still exhibit more bias than human-written references. Surprisingly, however, Mistral appears to be most susceptible to position biases and scores higher on extraction coverage than the smaller BART, T5, and Flan-T5 models. Overall, all models suffer from position biases and struggle to incorporate documents that appear later in inputs.
In addition to position biases, we examine how extraction coverage varies with length of the input documents in Figure 4. As longer documents are more likely to contain common terms, in this setting, we modify the extraction coverage metric to only consider terms unique to an individual input document. In doing so, we more precisely measure to what extent models extract from a particular document in relation to document length. In this case, we see a slight positive trend in extraction of unique terms with document length percentile among models compared to the human-written references [^11] For all models excluding Flan-T5, the extraction coverage of the top percentile of document lengths is significantly larger ($p\leq 0.05$) than that of the bottom percentile. We find that models tend to extract from longer documents, but the impact appears less pronounced than with document position.
Finally, we consider how the language of documents may influence models’ content selection. In particular, we consider how the density of highly arousing terms in an input document may lead it to be more likely to have information that will appear in a model’s summary, shown in Figure 5. In contrast to the position and length analyses, the relationship between extraction coverage and arousal of an input document is not monotonic. Instead, all models appear to extract less from both non-arousing and highly arousing documents, and to a lesser extent, reference summaries exhibit this behavior as well. This may be because models and humans struggle to incorporate underrepresented perspective outliers, that is, documents that present a perspective that diverges further from the other documents in the input. As with other analyses, the largest models appear to be most robust to this trend though still present a similar behavior. Therefore, contrary to our initial hypothesis, we conclude that models tend to avoid extracting from documents that use extremely many or few arousing terms.
@@ Appendix A Zero-Shot Baselines
Mistral, Mixtral, Vicuna, Llama-3, GPT-4o “Produce a short, single-sentence summary of the [left,right]-leaning political perspective within the following texts. Respond only with the summary beginning with "The [left,right]": {PASSAGES}”
@@ Appendix B Prompting Baselines
For larger LLMs, we use the second prompt in Table 7, generating each summary separately. Larger models (GPT-4o, Llama-3-70B, Vicuna-13B, and Mixtral-8x22B) are accessed via OpenAI and Fireworks AI [^12] https://fireworks.ai API’s. In all cases, we use a temperature of 0 to ensure deterministic outputs and greedy decoding. Smaller model (Mistral-7B, Llama-3-8B, and Vicuna-7B) inference is run with bfloat16 precision, also using greedy decoding.
@@ E.2 Results
Table 8: Average coverage and faithfulness scores for the Single-Perspective setting. Best performance on each metric is bolded and second best underlined. Higher scores reflect better performance for all metrics.
@@
Table 9: Perspective scores for all models in the Single-Perspective setting. S: Stance, O: Object, I: Intensity.
@@ F.3 Human Judgments by Perspective
S O I C R S O I C R
Ref. 2.9 3.1 2.8 2.9 1.7 2.9 3.4 2.9 2.9 2.0
Mistral 2.4 3.1 2.5 2.6 2.0 2.4 3.2 2.5 2.5 1.8
Mixtral 3.2 3.4 3.1 3.2 2.7 2.8 3.4 2.8 2.8 2.4
GPT-4o 3.3 3.5 3.3 3.3 2.9 3.2 3.6 3.2 3.2 2.8
Table 12: Average human judgments of each perspective dimension with average length and extraction coverage over all inputs. Scores are shown separately for left and right-perspective summaries.Best scores are bolded. Higher scores reflect better performance for all metrics. S: Stance, O: Object, I: Intensity, C: Consensus, R: Reasoning.
@@ F.4 Human Evaluation Interface
Screenshots of the human evaluation interface provided to annotators are shown in Figure 6, Figure 7, Figure 8, and Figure 9. Annotators are first provided with instructions describing each perspective dimension (Figure 6). Annotators are then shown example summaries with accompanying explanations describing how each perspective dimension could be interpreted in the example summary (Figure 7). We do not include input documents or possible judgments for the example summaries to avoid biasing annotators’ judgments. In the left-hand column, annotators are given the input passages separated by new lines, and in the right-hand column, annotators are provided with a summary followed by questions and 4-point Likert scales for each perspective dimension (Figure 8 & Figure 9). Reference summaries, and summaries generated by Mistral, Mixtral, and GPT-4o are shown on the same page in a randomized order for each set of input passages. Annotators are able to collapse and un-collapse all sections.
@@
Appendix G Extraction Analyses by Perspective
@@
Appendix H Additional Summary Examples

### v2

@@ 5.1 Qualitative Evaluation
Furthermore, Table 7 shows a selected pair of left and right summary examples where the right summary faithfulness metrics are notably larger than that of the left. In this case, neither summary contains noticeable faithfulness errors relative to the source documents, but there is a 2 point difference in SummaC scores and a 93 point difference in AlignScore both favoring the right-leaning summary. Throughout the dataset, cases resembling Table 7 where perspective summaries are faithful to similar extents support that neural metrics may suffer from political biases or spurious correlations beyond faithfulness of the input.
@@ 5.2 Extraction Analyses
To further understand model performance on this task, we analyze extraction behavior with respect to different qualities of the input documents: political perspective, position in the input, length, and use of arousing terms. We use extractive fragment coverage to quantify this behavior, as defined by Grusky et al. (2018). In analyses between the sets of left and right-leaning input documents, we hypothesize that models with better performance tend to extract more from the intended political perspective relative to the alternative perspective. Accordingly, we expect less performant models to extract similarly from both perspectives or more from the unintended perspective. For comparisons among documents sharing a common political perspective, we evaluate three additional hypotheses illustrating potentially undesirable behaviors: models extract more information from documents that (1) appear earlier in the input (position bias), (2) are longer (length bias), and (3) that contain more arousing language (arousal bias).
Document Perspective.
@@
As the Mixed-Perspective setting requires models to disentangle perspective among input documents, we first analyze whether models tend to extract more tokens from source documents with the correct political slant. To investigate this relationship, Figure 3 presents the extraction coverage scores for each combination of summary and source document perspective (e.g., left/right perspective summary and left/right-leaning source documents). For both perspectives, performant models tend to extract from source documents portraying the intended perspective (blue bars in the "Left Summaries" plot, and red bars in "Right Summaries") significantly more often than the alternative perspective. The less performant T5, Flan-T5, and Mistral models, however, present insignificant differences in the extent that they extract from the intended perspective with respect to the other perspective. In other words, these models extract similarly from left and right-leaning inputs regardless of the intended perspective of the summary. This suggests that poor performance of the models may be due to the inability to distinguish political leanings among input documents. Uniquely, BART consistently extracts significantly more from right-leaning documents than left-leaning, regardless of the intended perspective. Therefore, the largest and best performing models (e.g., GPT-4o, Llama-3.3, Llama-3.0, etc.) appear to be sufficiently capable of distinguishing perspectives in inputs.
@@
Document Position.
Position and lead biases in summarization systems are well-documented phenomena Chhabra et al. (2024); Olabisi and Agrawal (2024); Jung et al. (2019). Following these works, Figure 4 presents the extraction coverage scores across document positions for each model. As in prior research, we find a near monotonically decreasing relationship between extraction from a document and its position in the input, with all models drawing significantly more terms from the first document compared to those that follow. Among models, Llama-3.3 (70B) and GPT-4o appear to be most robust to document position, indicated by the flatter curves, but still exhibit more bias than human-written references. Surprisingly, however, Mistral appears to be most susceptible to position biases and scores higher on extraction coverage than the smaller BART, T5, and Flan-T5 models. Overall, all models suffer from position biases and struggle to incorporate documents that appear later in inputs.
Document Length.
In addition to position biases, we examine how extraction coverage varies with length of the input documents in Figure 4. As longer documents are more likely to contain terms shared with other documents, in this setting, we modify the extraction coverage metric to only consider terms unique to an individual input document. In doing so, we more precisely measure to what extent models extract from a particular document in relation to document length. In this case, we see a slight positive trend in extraction of unique terms with document length percentile among models compared to the human-written references [^12] For all models excluding Flan-T5, the extraction coverage of the top percentile of document lengths is significantly larger ($p\leq 0.05$) than that of the bottom percentile. We find that models tend to extract more unique terms from longer documents, but the impact appears less pronounced than with document position.
Document Arousal.
Finally, we consider how the language of documents may influence models’ content selection. In particular, we consider the relationship between the density of highly arousing terms in an input document and the extent to which the document is represented in the summary, shown in Figure 5. In contrast to the position and length analyses, the relationship between extraction coverage and arousal of an input document is not consistent across densities of arousing terms. Instead, all models appear to extract less from both non-arousing and highly arousing documents, and to a lesser extent, reference summaries exhibit this behavior as well. This may be because models and humans struggle to incorporate underrepresented perspectives (i.e., documents that present a perspective that diverges further from the other documents in the input). As with other analyses, the largest models appear to be most robust to this trend though still present a similar behavior. Therefore, contrary to our initial hypothesis, we conclude that models tend to avoid extracting from documents that use extremely many or extremely few arousing terms relative to other input documents.
@@ 9 Ethics Statement
Acknowledgments
This work was supported in part by grant IIS-2106666 from the National Science Foundation, the Knight First Amendment Institute at Columbia University, National Science Foundation Graduate Research Fellowship DGE-2036197, the Columbia University Provost Diversity Fellowship, and the Columbia School of Engineering and Applied Sciences Presidential Fellowship. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation or the Knight First Amendment Institute. We thank the anonymous reviewers and Smaranda Muresan for providing feedback on an earlier draft of the work.
@@ Appendix A Zero-Shot Baselines
Mistral, Mixtral, Vicuna, Llama-3.0, Llama-3.3, GPT-4o “Produce a short, single-sentence summary of the [left,right]-leaning political perspective within the following texts. Respond only with the summary beginning with "The [left,right]": {PASSAGES}”
@@ Appendix B Prompting Baselines
For larger LLMs, we use the second prompt in Table 8, generating each perspective summary separately. Larger models are accessed via OpenAI (GPT-4o) and Fireworks AI [^13] https://fireworks.ai (Llama-3.3-70B, Llama-3.0-70B, Vicuna-13B, and Mixtral-8x22B) API’s. In all cases, we use a temperature of 0 to ensure deterministic outputs and greedy decoding. Smaller model (Mistral-7B, Llama-3.0-8B, and Vicuna-7B) inference is run with bfloat16 precision, also using greedy decoding.
@@ E.2 Results
Table 9: Average coverage and faithfulness scores for the Single-Perspective setting. Best performance on each metric is bolded and second best underlined. For each model and metric, significantly higher left than right scores are highlighted in blue, and significantly higher right than left scores are highlighted in red. Higher scores reflect better performance for all metrics.
@@
Table 10: Perspective scores for all models in the Single-Perspective setting. Best performance on each metric is bolded and second best underlined. S: Stance, O: Object, I: Intensity.
@@ F.3 Human Judgments by Perspective
S O I C R Avg. Len Avg. Ext S O I C R Avg. Len Avg. Ext
Ref. 2.9 3.1 2.8 2.9 1.7 19.4 0.65 2.9 3.4 2.9 2.9 2.0 20.0 0.69
Mistral 2.4 3.1 2.5 2.6 2.0 47.1 0.83 2.4 3.2 2.5 2.5 1.8 50.0 0.84
Mixtral 3.2 3.4 3.1 3.2 2.7 45.7 0.71 2.8 3.4 2.8 2.8 2.4 46.8 0.73
GPT-4o 3.3 3.5 3.3 3.3 2.9 40.0 0.65 3.2 3.6 3.2 3.2 2.8 39.5 0.67
Table 13: Average human judgments of each perspective dimension with average length and extraction coverage over all inputs. Scores are shown separately for left and right-perspective summaries. Best scores are bolded. Higher scores reflect better performance for all metrics. Length is measured in Llama-3 tokens. S: Stance, O: Object, I: Intensity, C: Consensus, R: Reasoning.
@@ F.4 Human Evaluation Interface
Screenshots of the human evaluation interface using Label Studio Tkachenko et al. (2020-2024) provided to annotators are shown in Figure 6, Figure 7, Figure 8, and Figure 9. Annotators are first provided with instructions describing each perspective dimension (Figure 6). Annotators are then shown example summaries with accompanying explanations describing how each perspective dimension could be interpreted in the example summary (Figure 7). We do not include input documents or possible judgments for the example summaries to avoid biasing annotators’ judgments. In the left-hand column, annotators are given the input passages separated by new lines, and in the right-hand column, annotators are provided with a summary followed by questions and 4-point Likert scales for each perspective dimension (Figure 8 & Figure 9). Reference summaries, and summaries generated by Mistral, Mixtral, and GPT-4o are shown on the same page in a randomized order for each set of input passages. Annotators are able to collapse and un-collapse all sections.
@@
Appendix G Coverage and Faithfulness Metric Correlations
Appendix G presents correlations between human judgments on Mistral, Mixtral, and GPT-4o compared to each coverage and faithfulness metric. Among coverage metrics, Bleurt and BERTScore tend to be best correlated among left-perspective summaries, while Bleurt and ROUGE metrics tend to be best correlated among right-perspective summaries. The low correlation coefficients overall and particularly for faithfulness metrics, however, suggest that further investigation is needed to adapt coverage and faithfulness metrics to this domain.
Dimension Left Right
Coverage Faithfulness Coverage Faithfulness
R1 RL Bleurt BERT SummaC Align R1 RL Bleurt BERT SummaC Align
Stance -0.03 -0.06 0.22 0.14 0.06 0.19 0.10 0.12 0.24 0.00 0.02 0.15
Object 0.02 -0.01 0.25 0.02 0.08 0.13 0.10 0.13 0.18 0.11 0.07 -0.01
Intensity 0.05 -0.02 0.27 0.21 0.05 0.20 0.09 0.13 0.26 0.09 -0.01 0.12
Consensus 0.01 -0.03 0.24 0.12 0.01 0.18 0.17 0.18 0.27 0.02 -0.01 0.14
Reasoning -0.03 -0.06 0.20 0.14 -0.01 0.09 0.08 0.07 0.29 0.01 -0.00 0.15
Table 14: Kendall’s Tau correlations between the 5 perspective dimensions measured with human judgments, and the automatic coverage and faithfulness metrics. * indicates significant correlation ($p\leq 0.05$)
Appendix H Extraction Analyses by Perspective
@@
Appendix I Additional Summary Examples